OpenStudy (shameless):

what system of equations is represented by (-2,-2)and (5,5)? what system of equations is represented by (-5,3) and (-6,9) medal and fan for write answers.

OpenStudy (shameless):

Can someone plz help me

OpenStudy (shameless):

@AJ01

OpenStudy (anonymous):

ok

OpenStudy (shameless):

hey i really need help on this,u think you know it?

OpenStudy (anonymous):

okay you need to find the equation first

OpenStudy (shameless):

ok

OpenStudy (anonymous):

also the slope as

OpenStudy (anonymous):

what is the slope eq.

OpenStudy (shameless):

hey bro,i have the choices for them

OpenStudy (shameless):

if you need them

OpenStudy (anonymous):

the slope= (y2-y1)/x2-x1)

OpenStudy (shameless):

slope is y1-y2/x2-x1

OpenStudy (shameless):

yea lol

OpenStudy (anonymous):

okay find the slope

OpenStudy (shameless):

0

OpenStudy (anonymous):

y-yo=m(x-xo) pick one of the point

OpenStudy (shameless):

(5,5)?

OpenStudy (anonymous):

no...... the slope =(5-(-2))/(5-(-2)=7/7=1

OpenStudy (anonymous):

yeah.. but you slope is wrong

OpenStudy (anonymous):

*your

OpenStudy (shameless):

oh ok

OpenStudy (anonymous):

y-5=1(x-5)

OpenStudy (shameless):

aj I have the answer choices if you need them?

OpenStudy (anonymous):

no i don't

OpenStudy (shameless):

ok

OpenStudy (anonymous):

y-5=x-5 y=x-5+5 y=x

OpenStudy (shameless):

ok

OpenStudy (anonymous):

it is a linear equation

OpenStudy (shameless):

ik

OpenStudy (shameless):

but thats not one of the answers

OpenStudy (anonymous):

ok what are the answers

OpenStudy (anonymous):

straight line

OpenStudy (anonymous):

are you there !

OpenStudy (shameless):

this is for the first question a-y=-x-10/x-4 y=x b-y=x+10/x-4 y=-x c-y=-x+10/x-4 y=x d-y=-x+10/x-4 y=-x

OpenStudy (shameless):

sry it wasnt working

OpenStudy (anonymous):

is this the first part of the q

OpenStudy (shameless):

yea this is for the first qustion

OpenStudy (anonymous):

a) y=-x-(10/x) -4

OpenStudy (shameless):

how do you know?

OpenStudy (anonymous):

no just trying to rewrite the choices

OpenStudy (shameless):

oh ok

OpenStudy (shameless):

if i leave sometimes its because of my connection,sry bout that.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

m=1 y-y1=m(x-x1) y-5=(1)(x-5) y-5=x-5 y=x ----- y-(-2)=(1)(x-(-2)) y+2=x+2 y=x

OpenStudy (shameless):

ok